OpenStudy (anonymous):

Create a function, consisting of two different operations. Prove that your function is a legitimate function. Using your function, explain to how to solve for f(3). Show your work and explain each step using complete sentences. Also, explain how to find the inverse of your function. @Sheraz12345

OpenStudy (sheraz12345):

f(x) =3x+2 f(3)=3(3)+2=11 inverse wiill be y=3x+2 y-2=3x y-2/3=x so inverse is inverse= x-2/3

OpenStudy (anonymous):

Thankyou, but it must contain two different operations on each side of the equal sign.

OpenStudy (sheraz12345):

@ganeshie8 Plz look over to it ... by two operations she means y-2 =3x+2 ??????

OpenStudy (anonymous):

Noo, look this is an example of what she wants 2x + 4 = –3x + 14, this is from the text so I cannot use it.

OpenStudy (sheraz12345):

So make some other example like 3x+4=x-7

OpenStudy (anonymous):

That's good, so how do I prove it is an legitimate function?

OpenStudy (sheraz12345):

@ganeshie8 plZ

OpenStudy (sheraz12345):

To prove that certain equation is function or not we plug in different values of x and see if the answer for y is unique all the times .. But i just dont get it here

OpenStudy (sheraz12345):

@mathmate

OpenStudy (anonymous):

So, Instead of 3x+4=x-7, 3x+4=y-7

OpenStudy (sheraz12345):

yes it think so

OpenStudy (anonymous):

Thankyou c: